# adF: a Novel System for Measuring Web Fingerprinting through Ads

### v2

@@ 11.2. Fingerprinting countermeasures
The research community has been devoted to finding solutions to countermeasure the fingerprint Finding a definitive method or solution that avoids fingerprinting does not exist without compromising the user’s browsing experience. A widespread idea among the research community is to approach this issue to seek collective immunity, hide the user in the larger set of users, and try to standardize their attributes by modifying, overriding, randomizing, spoofing, etc. We find in the literature two types of proposals to counter fingerprinting.
@@
The first group proposes different types of manipulation of the reported attributes to third-parties (laperdrix2015mitigating, ; Al_FannahNasserMohammed2020Tltl, ). The solutions proposed by browser developers and introduced in Section 2 fall into this group. We also find proposals from the research community in this group. For instance, Blink (laperdrix2015mitigating, ) or Mimic (azad2020taming, ) propose to modify the content of the fingerprint by spoofing the value of certain attributes, whereas Canvas Blocker (canvasbloquer_firefox, ) blocks the whole Canvas. UniGL (wu2019rendered, ) proposes to universalize the fingerprint by removing discrepancies for the WebGL API between devices. The approach we propose in this paper, ShieldF, belongs to this group. Inspired by the work by Laperdrix et al. (laperdrix2016beauty, ), we do a thorough exploration of the attributes reported by browsers and mobile apps to third-parties to block those with a larger discrimination power, granted that they do not affect the end user’s browsing experience.
@@
The second group of solutions aims at identifying and blocking fingerprinting scripts. For instance, FP-Inspector (iqbal2021fingerprinting, ) uses machine learning classification algorithms and applies a layered set of restrictions to detect fingerprinting scripts. Privacy Badger (privacybadger, ) prevents the execution of known fingerprinting scripts in the browser. Also, a broader extension such as NoScript (noscript, ), which blocks the execution of unwanted JavaScript scripts, can be applied in the context of fingerprinting. This group of solutions are complementary to the one presented in this paper.
12. Conclusion
@@
Furthermore, we conduct an in-depth analysis of the discrimination power associated with different attributes reported by browsers and mobile apps. These findings hold significant value for developers of browsers and mobile apps seeking to implement more effective countermeasures against web fingerprinting by blocking or controlling those attributes with a higher discrimination power. Our detailed analysis also allows us to define a straightforward yet efficient solution to web fingerprinting. This solution proposes to block those attributes with the largest discrimination power, whose blocking does not affect the user experience in browsers or mobile apps. Our analysis shows that our proposal (easily adopted by browsers and mobile app developers) outperforms all the anti-fingerprinting solutions offered by major browser (Chrome, Safari and Firefox), reducing device vulnerability to web fingerprinting up to 62%. Given that collecting attributes from scripts embedded in ads may present more limitations compared to scripts embedded in web-pages’ HTML code, we believe that the results presented in this paper can be considered a lower bound of the vulnerability of devices to traditional web fingerprinting.
Appendix A Ethics considerations
The research and experiments described in this paper obtained the IRB approval of our institution through the Data Protection Officer, a member of the Ethics Committee responsible for approving projects and experiments with potential data protection implications.
Note that we take specific measures to prevent new risks for users (see Section 5). In particular, we only consider fingerprints for those users for which we have an advertising ID. By doing so, we avoid creating any unique ID for users who do not have one. Moreover, we do not process any data from devices that have the do-not-track flag active in their browser.
Finally, our adTag sends a total of 38,7kB to our back-end server. This is a negligible amount of data, which has a minimal impact on the overall data consumption of the device.
@@ Appendix A Attributes considered by the adF system
This appendix includes Table 4 providing detailed information of all the attributes initially considered to build the adF system.
Table 4. Attributes collected to build adF system. With asterisk (*) attributes considered for fingerprinting in the mobile app ecosystem.
@@
User Permissions state accelerometer - AJAX meth.
@@
WebGL Attributes antialias (quasi)-fixed value JS API
@@
WebGL ShaderPrecision: .rangeMin .rangeMax .precision -FRAGMENT (FR_S) -VERTEX (VR_S) FR_S_LOW_FLOAT correlated value JS API
@@
Audio formats .supported .smooth .powerEfficient acc [.supported]* - AJAX API

### v3

@@ XI-B Fingerprinting countermeasures
The research community has been devoted to finding solutions to countermeasure the fingerprint. Finding a definitive method or solution that avoids fingerprinting does not exist without compromising the user’s browsing experience. A widespread idea among the research community is to approach this issue to seek collective immunity, hide the user in the larger set of users, and try to standardize their attributes by overriding, randomizing, spoofing, etc. We find in the literature two types of proposals to counter fingerprinting.
@@
The first group proposes different types of manipulation of the reported attributes to third-parties [laperdrix2015mitigating, Al_FannahNasserMohammed2020Tltl]. The solutions proposed by browser developers and introduced in Section II fall into this group. We also find proposals from the research community. For instance, Blink [laperdrix2015mitigating] or Mimic [azad2020taming] propose to modify the content of the fingerprint by spoofing the value of certain attributes, whereas Canvas Blocker [canvasbloquer_firefox] blocks the whole Canvas. UniGL [wu2019rendered] proposes to universalize the fingerprint by removing discrepancies for the WebGL API between devices. The approach we propose in this paper, ShieldF, belongs to this group. Inspired by the work of Laperdrix et al. [laperdrix2016beauty], we do a thorough exploration of the attributes reported by browsers and mobile apps to third-parties to block those with a larger discrimination power, granted that they do not affect the end user’s browsing experience.
@@
The second group aims at identifying and blocking fingerprinting scripts. For instance, FP-Inspector [iqbal2021fingerprinting] uses machine learning classification algorithms and applies a layered set of restrictions to detect fingerprinting scripts. Privacy Badger [privacybadger] prevents the execution of known fingerprinting scripts in the browser. Also, a broader extension such as NoScript [noscript], which blocks the execution of unwanted JavaScript scripts, can be applied in the context of fingerprinting. This group of solutions are complementary to the one presented in this paper.
XII Conclusion
@@
Furthermore, we conduct an in-depth analysis of the discrimination power associated with different attributes reported by browsers and mobile apps. These findings hold significant value for developers of browsers and mobile apps seeking to implement more effective countermeasures against web fingerprinting by blocking or controlling those attributes with a higher discrimination power. Our detailed analysis also allows us to define a straightforward yet efficient solution to web fingerprinting. This solution proposes to block those attributes with the largest discrimination power, whose blocking does not affect the user experience in browsers or mobile apps. Our analysis shows that our proposal (easily adopted by browsers and mobile app developers) outperforms all anti-fingerprinting solutions offered by major browser (Chrome, Safari and Firefox), reducing device vulnerability to web fingerprinting up to 62%. Given that collecting attributes from scripts embedded in ads may present more limitations compared to scripts embedded in web-pages’ HTML code, we believe that results presented in this paper can be considered a lower bound of device vulnerability to traditional web fingerprinting.
Acknowledgement
We thank anonymous reviewers for their constructive feedback. This work is partially funded by framework of Recovery, Transformation and Resilience Plan funds, financed by European Union (Next Generation), under Grant Cátedras ENIA 2022 para la creación de cátedras universidad-empresa en IA with reference ETD/1180/2022. This paper is also partially carried out within framework of Recovery, Transformation and Resilience Plan funds, financed by European Union (Next Generation). Through grant Análisis y miTIgación de riesgos de seguridad y privaCIdad asociados a la explotación de datos PersonAles en OTTs (ANTICIPA). Further funding is allocated to ENTRUDIT, grant TED2021-130118B-I00 funded by MCIN/AEI/10.13039/5011000011033 and NextGeneration EU/PRTR funds.
Biography
[Uncaptioned image] Miguel A Bermejo-Agueda is a Ph.D. candidate in Telematics Engineering at Universidad Carlos III de Madrid. He received the B.Sc. in Telecommunications Engineering from Universidad Politécnica de Madrid and the M.Sc. in Data Science from the Universitat Oberta de Catalunya. His research interests include Internet measurements, online advertising, web sustainability, and privacy. He has published in venues such as ACM WWW and PoPETs, and has participated in EU H2020 projects.
[Uncaptioned image] Patricia Callejo is a Visiting Professor at Universidad Carlos III de Madrid, where she obtained her Ph.D. in Telematics Engineering (2020). She completed a research internship at ICSI, UC Berkeley, and received a RIPE Academic Cooperation Initiative (RACI) grant (RIPE 76, 2018). Her work spans Internet measurements, online advertising, and web transparency, with publications in ACM HotNets, ACM CoNEXT, and WWW, and participation in EU H2020 projects.
[Uncaptioned image] Rubén Cuevas is an Associate Professor at Universidad Carlos III de Madrid, Deputy Director of the UC3M–Santander Big Data Institute and Courtesy Assistant Professor, University of Oregon (2012). He has coauthored 70+ papers in venues including PNAS, CACM, IEEE/ACM TON, WWW, USENIX Security, ACM CoNEXT, and CHI. His interests include online advertising, web transparency, personalization and privacy, online social networks, and Internet measurements.
[Uncaptioned image] Ángel Cuevas is an Associate Professor at Universidad Carlos III de Madrid. He has coauthored 70+ papers in top venues, including PNAS, IEEE/ACM Transactions on Networking (TON), WWW, USENIX Security, CHI, and Communications of the ACM (CACM). His research focuses on Internet measurements, web transparency, privacy, and P2P networks. He received the Best Paper Award at ACM MSWiM 2010 and the 2018 Emilio Aced Research Prize from the Spanish Data Protection Agency.
[Uncaptioned image] Álvaro Mayol is Managing Director and Partner at TAPTAP Digital, previously serving as CPO and CTO. He is a Telecommunications Engineer from Universidad Politécnica de Madrid and the Illinois Institute of Technology. He began his career at Indra Systems (Electronic Warfare and Military Communications) and later led Solutions and Engineering at Geometry Global. In 2014 TAPTAP Digital acquired the company he co-founded, focused on proximity marketing and analytics.
@@ Appendix A Attributes considered by the adF system
Table IV provides detailed information of all the attributes initially considered to build the adF system and the mapping from individual attributes to meta-attributes. Low-level attributes that describe the same functional aspect (e.g., WebGL rendering parameters; navigator permissions) are concatenated into a single “meta-attribute”. While each constituent attribute may be only weakly discriminative on its own, jointly they capture a more stable, higher-level device property. Consolidating them removes redundancy and sparsity and yields a more compact representation—i.e., it effectively reduces dimensionality—while preserving the joint discriminative signal.
Table IV: Table IV: Attributes collected to build adF system. With asterisk (*) attributes considered for fingerprinting in the mobile app ecosystem.
@@
User Permissions
state
accelerometer - AJAX meth.
@@
WebGL
Attributes
antialias (quasi)-fixed value JS API
@@
WebGL
ShaderPrecision:
.rangeMin
.rangeMax
.precision
-FRAGMENT (FR_S)
-VERTEX (VR_S)
FR_S_LOW_FLOAT correlated value JS API
@@
Audio formats
.supported
.smooth
.powerEfficient
acc [.supported]* - AJAX API
@@
Appendix B Out-Of-Distribution (OOD) evaluation
To assess generalization beyond random, in-distribution validation, we evaluate the classifier under two out-of-distribution (OOD) settings: (i) cross-configuration (leave-one-configuration-out, LOCO) and (ii) temporal hold-outs across campaigns. In all cases, folds and splits are grouped by advertising ID (Ad-ID) so that all sessions from the same device remain on the same side of a split, preventing device-level leakage. The Ad-ID is used solely for ground-truth construction and grouping; it is never used as a model input.
Before any OOD experiment, we fix a single XGBoost configuration once, selected on a development split via 10-fold cross-validation grouped by Ad-ID. The very same hyperparameters are then reused for all OOD tests—no per-configuration re-tuning, no class reweighting, and no post-hoc threshold search—so the reported OOD figures constitute a conservative lower bound for a realistically pre-trained model.
To ensure reliability and avoid small-sample artefacts, we pre-specify inclusion thresholds independent of outcomes: (i) LOCO (cross-configuration): include {browser, OS} configurations with $\geq$ 2,000 unique Ad-IDs; (ii) Temporal hold-out: require $\geq$ 3 non-overlapping time windows with $\geq$ 3,000 unique Ad-IDs per window and disjoint Ad-IDs across time.
Due to configurations differ markedly in both class priors and test-set size, we report ROC-AUC (ranking, prior-agnostic) and balanced accuracy (thresholded, prior-sensitive), with 95% nonparametric bootstrap confidence intervals. To summarize across highly unbalanced test sizes (from $\approx$2.3k to $\approx$63k Ad-IDs), we provide three aggregates: macro (equal weight per configuration), size-weighted (by Ad-IDs/rows), and micro-pooled (concatenated test predictions).
Leave-One-Configuration-Out (LOCO)
we hold out one {browser, OS} configuration for testing and train on the rest (no Ad-ID overlap). Eight high-support configurations meet the inclusion criterion. Table V shows aggregated results: macro AUC $0.729$, macro balanced accuracy $0.579$; micro-pooled over all held-out tests: AUC $0.706$ [0.703, 0.709], balanced accuracy $0.652$ [0.650, 0.654]. As expected under sharp prior shifts—e.g., {Chrome, Windows} $\approx$96% positives vs. {Chrome, iOS} $\approx$4%—AUC transfers more stably across configs, whereas thresholded metrics reflect deployment
Table V: Table V: OOD evaluation. Values are point estimates with 95% bootstrap CIs where applicable.
Setting ROC-AUC Balanced accuracy
LOCO (macro) 0.729 0.579
LOCO (micro-pooled) 0.706  [0.703, 0.709] 0.652  [0.650, 0.654]
Temporal hold-out (2023–09/10) 0.769  [0.765, 0.773] 0.637  [0.634, 0.640]
priors (e.g., near-chance balanced accuracy in high-positive regimes despite AUC >0.62>0.62). This is precisely the scenario where simple per-config calibration (e.g., Platt, isotonic) on the training portion aligns decision thresholds without changing the model.
Temporal hold-out.
We train on the first three predefined windows (02-2022, 05/06-2022, 05-2023) and test prospectively on 09/10-2023. With the same fixed hyperparameters, the model attains AUC $0.769$ [0.765, 0.773] and balanced accuracy $0.637$ [0.634, 0.640] on the held-out window (see Table V), indicating temporal generalization under natural {browser, OS} drift.
OOD results complement the in-distribution, grouped cross-validation reported in the main text and extend the reach of the ground-truth statistic ($\mathcal{TV}$) by providing deployable, identifier-free risk estimates on unseen configurations and future windows. Residual, prior-driven threshold mismatch is expected and can be addressed via light-weight calibration at deployment time; we intentionally keep the reported OOD figures uncalibrated to provide a conservative lower bound.
Appendix C Temporal Stability of Configuration-Level Web Fingerprinting
To complement our core evaluation and address potential concerns regarding the temporal robustness of fingerprinting, we conduct a longitudinal analysis across four timeframes in our dataset: 02-2022, 05/06-2022, 05-2023, and 09/10-2023. Unlike prior works as Laperdrix et al. (2020) that examine the persistence of individual fingerprints over sessions, our approach focuses on the temporal evolution of the fingerprinting landscape at the configuration level, defined as the tuple {browser, OS}, as observed through repeated ad delivery campaigns.
While Section IX.C analyzes the individual attribute entropy, here we extend the perspective using the joint entropy of all fingerprinting attributes to capture the overall structural complexity of the fingerprint space over time. Formally, the Shannon joint entropy of a set of attributes $X_{1},X_{2},\ldots,X_{n}$ is:
$H(X_{1},X_{2},...,X_{n})=-\sum_{(x_{1},x_{2},...,x_{n})\in\mathcal{X}}P(x_{1},x_{2},...,x_{n})\\ \times\log_{2}P(x_{1},x_{2},...,x_{n})$ (4)
where $\mathcal{X}$ is the set of all possible combinations of attribute values, and $P(x_{1},x_{2},...,x_{n})$ is the joint empirical probability of each such combination. This measure captures the richness and dispersion of fingerprints independently of class imbalance or sample size.
To ensure reliability, we only include configurations that (i) maintain at least 400 samples per timeframe, and (ii) exhibit a sample count variability (standard deviation over mean) below 70% across timeframes. These constraints reduce bias from underrepresented or unstable datasets. The resulting set includes configurations for iOS and macOS, which meet these criteria consistently. Figure 6 summarizes their temporal evolution.
Figure 6: Joint entropy ($\mathcal{H}$) over time for selected browser–OS configurations meeting sample stability criteria. Desktop setups show consistently higher entropy, while iOS-based configurations maintain lower and stable values.
Among desktop setups, {Chrome, macOS} maintains high joint entropy (11.3–12.2 bits), indicating a persistent and diverse fingerprint space—suggesting little evolution in defenses. In contrast, {Safari, macOS} remains more stable and less expressive (10.0–11.0 bits), consistent with reduced API exposure. On mobile, {Safari, iOS} shows a notably stable and low joint entropy (9.9–11.8 bits), reaffirming its well-known fingerprinting resistance. {Chrome, iOS} exhibits the lowest entropy values across all configurations (<<10.3 bits), likely due to iOS sandboxing that limits Chrome’s attribute exposure. This behavior contrasts with desktop counterparts and supports previous findings that attribute diversity on iOS is intentionally constrained by platform-level controls.
No major non-monotonic shifts are observed among the selected configurations. Minor fluctuations (e.g., a dip for {Chrome, iOS} in 05 - 2023) correspond to lower sample sizes and should be interpreted with caution. These trends align with the findings of Laperdrix et al. (2020), who report that fingerprint changes are often driven by changes in browser APIs, software updates, and hardware context—not necessarily by user behaviour or population shifts.
In summary, fingerprinting vulnerability does evolve over time, albeit unevenly across platforms and browsers. Joint entropy emerges as a robust metric to capture these temporal dynamics and validates the ability of adF to operate as a longitudinal and configuration-aware fingerprinting auditor under real deployment conditions.
Appendix D Ethics considerations
The research and experiments described in this paper obtained the IRB approval of our institution through the Data Protection Officer (DPO), a member of the Ethics Committee responsible for approving projects and experiments with potential data protection implications. Note that we take specific measures to prevent new risks for users (see Section V). In particular, we only consider fingerprints for those users for which we have an advertising ID. By doing so, we avoid creating any unique ID for users who do not have one. Moreover, we do not process any data from devices that have the do-not-track flag active in their browser. Importantly, Advertising IDs are never used as features by our classifier. They are used exclusively to (i) derive the ground-truth device identity for evaluation, and (ii) define non-leaking train/test splits via grouping.
Finally, our adTag sends a total of 38.7kB per measurement to our back-end server. This is a negligible amount of data, which has a minimal impact on the overall data consumption of the device.